# Calibration of a $\Delta$E-E telescope based on CeBr3 scintillator for secondary charged particles measurements in hadron therapy

### v1

@@ 3.3 Telescope ions identification
The identification of different ions using the $\Delta E$-$E$ telescope leverages the energy deposition characteristics in the plastic and CeBr3 scintillators. Figure 8 demonstrates the ability of the $\Delta E$-$E$ telescope to distinguish ions based on their charge state (Z). The plot shows the energy deposited in the plastic scintillator ($\Delta E$) as a function of the energy deposited in the CeBr3 scintillator ($E$), for the secondary particles produced by a 200 MeV/u ${}^{12}C$ ion beam on a RW3 target (tissue equivalent).
Each ion species (Z = 1 to Z = 6) forms a distinct branch in the $\Delta E$-$E$ space, reflecting its unique energy loss and remaining energy characteristics. The ability to resolve these branches accurately enables the identification of secondary charged particles generated during hadrontherapy scenario.
@@ 3.4 Time performance
Figure 9: Resolution in time between the plastic with the CeBr3 and R7057 PMT with a 25MeV 1H beam (a) and XP3990 PMT with a 180 MeV/u 12C beam (b)
The temporal resolution obtained from those distributions for the 180 MeV/u 12C beam, is of 282 $\pm$ 1 ps. Similarly, for the plastic and CeBr3 detectors with 25 MeV protons beam, the time resolution was determined of 171 $\pm$ 1 ps. The increase of the time resolution value for carbon ions is mainly due to the low voltage that was applied to the cerium bromide PMT (+ 350 V). These values represent crucial benchmarks in assessing the time performance of the detection system across varying beam energies and particle types. Such precise temporal resolution enables robust particle identification and TOF measurement, and enhances the overall capabilities of the experimental setup for the CLINM purpose.
@@ 4 Discussion
In addition to energy resolution, the time performance of our $\Delta$E-E telescope was also evaluated. Precise time measurements enable the correlation of energy loss and total energy, facilitating particle identification and background rejection. Unlike prior systems optimized for primary beam monitoring, this setup is uniquely calibrated for the detection of secondary charged particles, addressing a critical gap in data availability for hadrontherapy. The time resolution measurements exhibit a resolution of 282 $\pm$ 1ps for 12C ions of 180 MeV/u and 171 $\pm$ 1ps for 25 MeV protons. These results underscore the capability of this telescope to provide precise measurements in clinical conditions. The narrower time resolution for protons compared to 12C ions can be attributed to the differences in their masses and velocities. The timing resolution of 171–282 ps, coupled with an energy resolution on the order of 10 MeV, represents a substantial advancement in the detection of secondary particles. This precision enables not only better particle identification but also provides accurate input for Monte Carlo models, thereby improving the reliability of dose calculations.
@@ 5 Conclusion
In the context of hadrontherapy, precision in measuring energy deposition stands as a fundamental requirement for optimizing treatment plans.
In this work, the calibration and performance evaluation of a $\Delta$E-E telescope was presented, designed for the detection of secondary charged particles generated in hadrontherapy under clinically relevant conditions. This telescope combines a thin plastic scintillator and a CeBr3 crystal scintillator, optimized for precision in energy and time measurements.
@@
The energy resolution was determined to be on the order of 10 MeV for both scintillators, ensuring precise identification of secondary particles. The system’s time resolution was measured at 282 $\pm$ 1 ps for 180 MeV/u 12C ions and 171 $\pm$ 1 ps for 25 MeV protons. These results position the telescope as a robust tool for $\Delta$E-E and $\Delta$E-ToF measurements, capable of separating secondary particles generated during beam-tissue interactions.
In clinical terms, the telescope operates effectively within the energy ranges and ion types encountered in hadrontherapy. By accurately characterizing secondary charged particles, this study addresses a critical gap in experimental data, providing valuable input for Monte Carlo simulations and improving treatment planning accuracy.

### v2

@@ 3.3 Telescope ions identification
The identification of different ions using the $\Delta E-E$ telescope leverages the energy deposition characteristics in the plastic and CeBr3 scintillators. Figure 8 demonstrates the ability of the $\Delta E-E$ telescope to distinguish ions based on their charge state (Z). The plot shows the energy deposited in the plastic scintillator ($\Delta E$) as a function of the energy deposited in the CeBr3 scintillator ($E$), for the secondary particles produced by a 200 MeV/u ${}^{12}C$ ion beam on a RW3 target (tissue equivalent), revealing distinct structures. This figure is intended as a qualitative illustration of the telescope’s ability to discriminate charged fragments by charge (Z) using $\Delta E-E$ correlations. A more detailed analysis, including Monte Carlo comparison, separation efficiency, and species contamination rates, is beyond the scope of this work and will be presented in a dedicated publication.
Each ion species (Z = 1 to Z = 6) forms a distinct branch in the $\Delta E-E$ space, reflecting its unique energy loss and remaining energy characteristics. The contours in the Figure are intended as visual guides based directly on the observed data, to emphasize visible patterns in the distributions. This demonstrates the ability of the $\Delta E-E$ telescope to identify secondary charged particles produced by therapeutic ion beams interacting with tissue-equivalent targets under experimental conditions representative of clinical hadrontherapy.
@@ 3.4 Time performance
Figure 9: Resolution in time (measured time difference) between the plastic with the CeBr3 and R7057 PMT with a 25 MeV 1H beam (a) and XP3990 PMT with a 180 MeV/u 12C beam (b)
The temporal resolution obtained from those distributions for the 180 MeV/u 12C beam, is of 282 $\pm$ 1 ps. Similarly, a time resolution of 171 $\pm$ 1 ps was measured between the plastic and CeBr3 detectors using a 25 MeV proton beam.
The uncertainties on the time resolution correspond to statistical errors only. The increase of the time resolution value for carbon ions is mainly due to the low voltage that was applied to the cerium bromide PMT (+ 350 V). These values represent crucial benchmarks in assessing the time performance of the detection system across varying beam energies and particle types. Such temporal resolution demonstrates the system’s potential for future implementation of $\Delta E$–ToF particle identification methods, and enhances the overall capabilities of the experimental setup for the CLINM purpose.
@@ 4 Discussion
In addition to energy resolution, the time performance of our $\Delta E-E$ telescope was also evaluated. Precise time measurements enable the correlation of energy loss and total energy, facilitating particle identification and background rejection. By precisely calibrating and characterizing the $\Delta E-E$ telescope, this study establishes a robust foundation for future measurements of secondary charged particles. Such data, once acquired in tissue-equivalent material, will provide valuable input for Monte Carlo benchmarking and, in the long term, may help improve the accuracy of treatment planning systems.
The time resolution measurements exhibit a resolution of 282 $\pm$ 1 ps for 12C ions of 180 MeV/u and 171 $\pm$ 1 ps for 25 MeV protons. These results underscore the capability of this telescope to provide precise measurements under experimental conditions that are representative of clinical hadrontherapy scenarios. The narrower time resolution distribution for protons compared to 12C ions can be attributed to the differences in their masses and velocities. The timing resolution of 171–282 ps, coupled with an energy resolution on the order of 10 MeV, represents a substantial advancement in the detection of secondary particles. This precision enables not only better particle identification but also provides accurate input for Monte Carlo models, thereby improving the reliability of dose calculations.
@@
It should be emphasized again that Birks’ law was used here as an effective parametrization of the system response. The extracted constants are not intrinsic to the CeBr3 material itself, but reflect the global response of the detector system under specific operating conditions. This empirical modeling approach is particularly suited to experimental configurations where full analytical modeling of light yield and collection is impractical.
@@ 5 Conclusion
In the context of hadrontherapy, precision in measuring energy deposition stands as a fundamental requirement for optimizing treatment plans. In this work, the calibration and performance evaluation of a $\Delta E-E$ telescope was presented, designed for the detection of secondary charged particles generated by ion beams under experimental conditions representative of clinical hadrontherapy. This telescope combines a thin plastic scintillator and a CeBr3 crystal scintillator, optimized for precision in energy and time measurements.
@@
The energy resolution was determined to be on the order of 10 MeV for both scintillators, ensuring precise identification of secondary particles. The system’s time resolution was measured at 282 $\pm$ 1 ps for 180 MeV/u 12C ions and 171 $\pm$ 1 ps for 25 MeV protons. These results confirm the potential of the telescope as a reliable system for $\Delta E-E$ and $\Delta E$-ToF measurements, capable of separating secondary particles generated during beam-tissue interactions.
The telescope was tested under experimental conditions designed to replicate the physical characteristics of clinical hadrontherapy, and was shown to perform effectively within the relevant energy ranges and ion species. The system is not intended for direct use on patients, but rather for producing precise and reliable nuclear data using tissue-equivalent targets. These measurements provide essential input for Monte Carlo simulations, ultimately improving the accuracy of treatment planning systems.